# Transformer-Based Fault-Tolerant Control for Fixed-Wing UAVs Using Knowledge Distillation and In-Context Adaptation

### v1

@@ Abstract
This study presents a transformer-based approach for fault-tolerant control in fixed-wing Unmanned Aerial Vehicles (UAVs), designed to adapt in real time to dynamic changes caused by structural damage or actuator failures. Unlike traditional Flight Control Systems (FCSs) that rely on classical control theory and struggle under severe alterations in dynamics, our method directly maps outer-loop reference values—altitude, heading, and airspeed—into control commands using the in-context learning and attention mechanisms of transformers, thus bypassing inner-loop controllers and fault-detection layers. Employing a teacher-student knowledge distillation framework, the proposed approach trains a student agent with partial observations by transferring knowledge from a privileged expert agent with full observability, enabling robust performance across diverse failure scenarios. Experimental results demonstrate that our transformer-based controller outperforms industry-standard FCS and state-of-the-art reinforcement learning (RL) methods, maintaining high tracking accuracy and stability in nominal conditions and extreme failure cases, highlighting its potential for enhancing UAV operational safety and reliability.
@@ I Introduction
Unlike other robotics fields, such as manipulation and humanoid locomotion, where advanced control methods are essential for managing complex joint movements, UAV Flight Control Systems (FCSs) in industry typically rely on classical control techniques for low-level control layers. While modern approaches, like Model Predictive Control (MPC), offer significant advantages for high-level tasks such as trajectory planning and collision avoidance [1, 2], they require precise system models, extensive uncertainty handling, and high computational resources, which often make them impractical for low-level UAV control. The simplicity, reliability, and efficiency of classical control techniques have established them as the preferred choice for UAVs attitude control.
Figure 1: Trajectory comparison between the proposed transformer-based controller (blue) and an industry-standard FCS (red). Figures (a) and (b) illustrate nominal scenario tracking, while (c) and (d) demonstrate the controllers’ responses to semi-wing damage, with the FCS losing control and the proposed method stabilizing the UAV.
@@
Fault-tolerant flight control has become a focal point for safety-critical UAV operations. Typically, fault-tolerant methods rely on fault detection and diagnosis techniques, which identify faults and then adjust controller parameters to account for the new dynamics [3, 4, 5, 6]. This approach is complex, requiring real-time fault identification and parameter adjustment in a highly nonlinear dynamic environment.
@@
Reinforcement Learning (RL) has introduced alternative solutions to the problem [7, 8, 9, 10]. With its ability to handle high-dimensional, nonlinear dynamics, RL holds promise for system fault management. However, RL algorithms are typically designed for Markov Decision Process (MDP) formulations, while fault-tolerant control—with sudden, unobserved changes in dynamics—must be framed as a Partially Observable Markov Decision Process (POMDP), making RL algorithms learning more difficult and possibly leading them to suboptimal performance [11]. Although recent works have shown possible solutions to this problem [11, 12, 13], these add considerable complexity to the algorithms. Most research focuses on actuator faults in multicopter UAVs [4, 9] or on fixed-wing UAV fault tolerance at the inner attitude control loop level, avoiding altitude, heading, and airspeed tracking [7, 8].
@@
In this work, we propose a novel transformer-based fault-tolerant control method to directly map reference points in the outer loop of attitude UAV control—altitude ($h$), heading ($\Psi$), and airspeed ($V_{T}$)—into control-surface and throttle commands, enabling full UAV control without the need for inner control loops, which complicate the system. Our method employs the attention mechanisms and in-context learning capabilities of transformer models to adapt control actions to dynamic changes during inference, thereby eliminating the need for fault detection or identification and parameter adjustments in the Flight Control System (FCS). Our transformer-based controller uses a context window of past UAV states to autonomously detect and adapt to dynamic changes. Fig. 1 shows a comparison between our proposed transformer-based controller (blue trajectory) and an industry-standard FCS (red trajectory). In Fig. 1(a) and Fig. 1(b), both systems track the commanded references in nominal conditions. Figs. 1(c) and 1(d) illustrate the response when the UAV experiences semi-wing damage, showing the FCS losing control and causing a crash, while our method stabilizes the UAV and follows the reference values.
@@
  • We design a transformer-based flight controller that utilizes in-context learning to adapt its behavior in real-time as UAV dynamics change due to failures. This controller employs a context of past states to determine actions, eliminating the need for fault detection methods.
@@ II-B Transformers
Transformers, known for their attention mechanism and advantages in sequence modeling, were initially applied in the field of Natural Language Processing (NLP) [17]. Owing to their high-quality global contextual learning and efficient parallel computation, the transformer architecture has emerged as a powerful alternative to traditional sequence prediction methods like Recurrent Neural Networks (RNNs) and Long Short-Term Memory Networks (LSTMs).
With advances in transformers, their application has extended to RL [18, 19]. Researchers have employed transformer architectures in decision-making agents, proposing the Decision Transformer (DT) model, which frames the sequence modeling problem as an action prediction problem based on historical states and reward-to-go sequences [18]. Trained in a supervised fashion, DT seeks to minimize the error between predicted and ground-truth actions by using sequences of past states and reward-to-go. Recent work has shown that DTs can outperform many state-of-the-art offline RL algorithms [18]. In this work, a DT model is trained to serve as an attitude controller for a UAV in failure scenarios, outperforming other state-of-the-art methods.
@@ III Problem Statement
These equations define the aerodynamic coefficients $C_{L}$, $C_{D}$, $C_{Y}$, $C_{l}$, $C_{m}$, and $C_{n}$, which represent the lift, drag, side-force, roll moment, pitch moment, and yaw moment coefficients, respectively [26]. Each equation is influenced by the UAV’s dynamic variables (e.g., angle of attack $\alpha$, sideslip angle $\beta$, pitch rate $q$, roll rate $p$, yaw rate $r$) and control surface deflections (e.g., elevator $\delta_{e}$, aileron $\delta_{a}$, rudder $\delta_{r}$). In these expressions, $c$ refers to the mean aerodynamic chord, $b$ is the wingspan of the UAV, $k$ is a factor related to induced drag, and $V$ is the UAV’s airspeed. The coefficients are functions of the stability and control derivatives, which are significantly affected by the UAV’s aerodynamic configuration and can be substantially altered in cases of faults or damage.
@@ IV-A Privileged Agent Training through Online Reinforcement Learning
The privileged agent is trained using the DreamerV3 algorithm [12], selected for its efficiency, robustness, and ability to generalize across multiple environments without extensive hyperparameter tuning. DreamerV3 is an online model-based RL algorithm that learns a world model of the environment from past trajectories, then refines a policy through imagination across the latent space learned in the world model encoder.
@@ V-A Testing Cases and Implementation Details
We evaluate our method across several failure and damage scenarios for the UAV. Unlike the random and independent parameter modifications used during training, testing cases involve modifications to aerodynamic parameters and control surfaces based on the changes that a specific failure produces in the UAV’s dynamics, creating unseen situations for the agent. The defined scenarios are:
@@
Experiments and training are conducted using JSBSim, a high-fidelity 6-DoF flight dynamics model. The UAV physics model used operates within flight conditions of $h\in[4000,24000]$ ft and $V_{T}\in[260,360]$ ft/s. Episodes have a maximum length of 1000 steps. Light-intensity wind gusts are enabled during validation, following a Gaussian distribution N(0, 2 ft/s) for the north, east, and down gust components.
@@ V-B Comparison Study
Damaged Wing $\mu\pm\sigma$310.98 ± 23.46 -349.55 ± 71.68 -1394.93 ± 155.40 -1167.78 ± 338.98 -762.48 ± 444.57
Crash % 0.00% 0.00% 100.00% 92.00% 60.00%
@@
To assess the effectiveness of the DT under POMDP conditions, we compare the results of the DT student agent and the privileged teacher agent. Both policies are tested across all scenarios in 100 episodes each, with flight conditions and random seeds standardized for fair comparison. Fig. 4 shows the results, using the mean episode return as the performance metric, where the theoretical maximum is zero. Results demonstrate similar performance between the agents, showing the DT’s capability to learn behaviors from the expert dataset under the POMDP setting and to adapt to dynamic variations in context. In some scenarios, the DT even outperforms the teacher agent, indicating that it learns through offline RL rather than simply mimicking the training data through imitation learning (IL).
@@
Our method takes advantage of the in-context learning capabilities of transformers to adapt policy behavior during flight without changing weights, rather than modifying controller gains. Fig. 5 presents sample trajectories of our method versus the other algorithms in the comparison study for tracking $h$, $\Psi$, and $V_{T}$. In Fig. 5(a), the nominal UAV dynamics are shown, where all designed controllers track reference values with similarly high performance, with some noise in the measured airspeed due to light wind gusts.

### v2

@@ Abstract
This study presents a transformer-based approach for fault-tolerant control in fixed-wing Unmanned Aerial Vehicles (UAVs), designed to adapt in real time to dynamic changes caused by structural damage or actuator failures. Unlike traditional Flight Control Systems (FCSs) that rely on classical control theory and struggle under severe alterations in dynamics, our method directly maps outer-loop reference values—altitude, heading, and airspeed—into control commands using the in-context learning and attention mechanisms of transformers, thus bypassing inner-loop controllers and fault-detection layers. Employing a teacher-student knowledge distillation framework, the proposed approach trains a student agent with partial observations by transferring knowledge from a privileged expert agent with full observability, enabling robust performance across diverse failure scenarios. Experimental results demonstrate that our transformer-based controller outperforms industry-standard FCS and state-of-the-art reinforcement learning (RL) methods, maintaining high tracking accuracy and stability in nominal conditions and extreme failure cases, highlighting its potential for enhancing UAV operational safety and reliability. Video: https://youtu.be/ATW3LZFRqc0
Index Terms:
Aerial Robotics, Learning-based Control, Reinforcement Learning, Fault-Tolerant Control.
@@ I Introduction
Figure 1: Trajectory comparison between the proposed transformer-based controller (blue) and an industry-standard FCS (red). Figures (a) and (b) illustrate nominal scenario tracking, while (c) and (d) demonstrate the controllers’ responses to semi-wing damage, with the FCS losing control and the proposed method stabilizing the UAV
In contrast to robotics fields like manipulation and humanoid locomotion—which rely on advanced control methods to manage complex joint movements—UAV flight control systems (FCSs) in industry typically use classical control techniques for low-level control. Although modern approaches such as Model Predictive Control (MPC) offer significant advantages for high-level tasks like trajectory planning and collision avoidance [1, 2], they require precise system models, extensive uncertainty handling, and high computational resources, which often makes them impractical for low-level UAV control. Instead, the simplicity, reliability, and efficiency of classical control techniques—such as cascade Proportional-Integral-Derivative (PID) control systems—have made them the preferred choice for UAV attitude control, and they are implemented in widely deployed autopilots like PX4 and Ardupilot.
@@
Fault-tolerant flight control has become a focal point for safety-critical UAV operations. Typically, fault-tolerant methods rely on fault detection and diagnosis techniques, which identify faults and then adjust controller parameters to account for the new dynamics [3, 4, 5, 6]. This approach is complex, requiring real-time fault identification and parameter adjustment in a highly nonlinear dynamic environment. Other studies have explored novel approaches to nonlinear fault-tolerant control without explicit fault identification. These include approaches based on fuzzy logic systems [7], sliding mode control [8], and robust control techniques [9].
@@
Reinforcement Learning (RL) has introduced alternative solutions to the problem [10, 11, 12, 13]. With its ability to handle high-dimensional, nonlinear dynamics [14], RL holds promise for system fault management. However, RL algorithms are typically designed for Markov Decision Process (MDP) formulations, while fault-tolerant control—with sudden, unobserved changes in dynamics—must be framed as a Partially Observable Markov Decision Process (POMDP), making RL algorithms learning more difficult and possibly leading them to suboptimal performance [15]. Although recent works have shown possible solutions to this problem [15, 16, 17], these add considerable complexity to the algorithms. Most research focuses on actuator faults in multicopter UAVs [4, 12] or on fixed-wing UAV fault tolerance at the inner attitude control loop level, avoiding altitude, heading, and airspeed tracking [10, 11].
Although learning-based control methods like Reinforcement Learning (RL) and Imitation Learning (IL) have great potential, ensuring stability with these techniques is more complex than with traditional control theory algorithms due to their intrinsic non-linearity and the lack of a model of the system. Several studies have attempted to integrate stability guarantees into RL and IL-based algorithms [18, 19]. Other works, based on Lyapunov theory, focus on learning a candidate Lyapunov function–commonly referred to as a Neural Lyapunov function [20, 21, 22, 23]. In Ref. [24], instead of simply penalizing violations of the Lyapunov condition without labeled data for direct regression, the authors utilize numerical tools based on Koopman Operator theory to generate explicit ground truth data for training.
@@
In this work, we propose a novel transformer-based fault-tolerant control method to directly map reference points in the outer loop of attitude UAV control—altitude ($h$), heading ($\Psi$), and airspeed ($V_{T}$)—into control-surface and throttle commands, enabling full UAV control without the need for inner control loops, which complicate the system. Our method employs the attention mechanisms and in-context learning capabilities of transformer models to adapt control actions to dynamic changes during inference, thereby eliminating the need for fault detection or identification and parameter adjustments in the Flight Control System (FCS). Our transformer-based controller uses a context window of past UAV states to autonomously detect and adapt to dynamic changes, similar to the ideas described in Ref. [25]. Moreover, we incorporate Lyapunov-based concepts into the offline RL training process to encourage stable model behavior. Fig. 1 shows a comparison between our proposed transformer-based controller (blue trajectory) and an industry-standard FCS (red trajectory). In Fig. 1(a) and Fig. 1(b), both systems track the commanded references in nominal conditions. Figs. 1(c) and 1(d) illustrate the response when the UAV experiences semi-wing damage, showing the FCS losing control and causing a crash, while our method stabilizes the UAV and follows the reference values.
@@
  • We design a transformer-based flight controller that utilizes in-context learning to adapt its behavior in real-time as UAV dynamics change due to failures. This controller employs a context of past states to determine actions, eliminating the need for fault detection methods. Moreover, Lyapunov-based concepts are included to encourage controller stability.
@@ II-B Transformers
Transformers, known for their attention mechanism and advantages in sequence modeling, were initially applied in the field of Natural Language Processing (NLP) [29]. Owing to their high-quality global contextual learning and efficient parallel computation, the transformer architecture has emerged as a powerful alternative to traditional sequence prediction methods like Recurrent Neural Networks (RNNs) and Long Short-Term Memory Networks (LSTMs). From a robotics and control point of view, transformers outperform recurrent models in capturing long-range dependencies and rapidly adapting to dynamic system variations, as demonstrated in Ref. [30]. Later in this work, a performance comparison between these models is conducted.
With advances in transformers, their application has extended to RL [31, 32, 33]. Researchers have employed transformer architectures in decision-making agents, proposing the Decision Transformer (DT) model, which frames the sequence modeling problem as an action prediction problem based on historical states and reward-to-go sequences [31]. Trained in a supervised fashion, DT seeks to minimize the error between predicted and ground-truth actions by using sequences of past states and reward-to-go. Recent work has shown that DTs can outperform many state-of-the-art offline RL algorithms [31]. In this work, a DT model is trained to serve as an attitude controller for a UAV in failure scenarios, outperforming other state-of-the-art methods.
@@ III Problem Statement
These equations define the aerodynamic coefficients $C_{L}$, $C_{D}$, $C_{Y}$, $C_{l}$, $C_{m}$, and $C_{n}$, which represent the lift, drag, side-force, roll moment, pitch moment, and yaw moment coefficients, respectively [38, 39]. Each equation is influenced by the UAV’s dynamic variables (e.g., angle of attack $\alpha$, sideslip angle $\beta$, pitch rate $q$, roll rate $p$, yaw rate $r$) and control surface deflections (e.g., elevator $\delta_{e}$, aileron $\delta_{a}$, rudder $\delta_{r}$). In these expressions, $c$ refers to the mean aerodynamic chord, $b$ is the wingspan of the UAV, $k$ is a factor related to induced drag, and $V$ is the UAV’s airspeed. The coefficients are functions of the stability and control derivatives, which are significantly affected by the UAV’s aerodynamic configuration and can be substantially altered in cases of faults or damage.
@@ IV-A Privileged Agent Training through Online Reinforcement Learning
The privileged agent is trained using the DreamerV3 algorithm [16], selected for its efficiency, robustness, and ability to generalize across multiple environments without extensive hyperparameter tuning. DreamerV3 is an online model-based RL algorithm that learns a world model of the environment from past trajectories, then refines a policy through imagination across the latent space learned in the world model encoder.  [^1] After publication, the code will be available at: https://github.com/fgiral000/ft-decision-transformer
@@ IV-B Knowledge Distillation Via Decision Transformer
IV-C Neural Lyapunov Function Learning
To encourage stable behavior of the control policy, we incorporate a Neural Lyapunov function $V_{\theta}(s)$ into our training process, following ideas developed at Refs. [21, 22, 24]. Inspired by [24], we supervise the Lyapunov function using Koopman eigenfunctions computed on the fly from trajectory data.
Taking advantage of the trajectories dataset generated by the privileged agent, and considering as the Lyapunov state $s\in[\epsilon_{h},\epsilon_{\Psi},\epsilon_{V_{T}},\phi,p,q,r]$, we first define an observable $f(s)$, which we choose as the L2 norm of the state:
$f(s)=\|s\|_{2}.$ (12)
The Koopman eigenfunction associated with the initial state $s_{0}$ is then approximated using a Laplace average along the trajectory:
$\phi(s_{0})\approx\frac{\sum_{t=0}^{T-1}e^{-\lambda t}\,f(s_{t})}{\sum_{t=0}^{T-1}e^{-\lambda t}},$ (13)
where $\lambda>0$ is a decay parameter and $T$ is the sequence length limited to 60 steps. This weighted average provides a measure that decays along trajectories, capturing the contraction properties required of a Lyapunov function.
The Neural Lyapunov function, parameterized by a multi-layer perceptron (MLP), is trained to approximate the Koopman eigenfunction values. The MLP consists of two hidden layers with 256 neurons each, followed by ReLU activations, and an output layer with a squared activation function. The regression loss is defined as the mean squared error (MSE):
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\left(V_{\theta}(s_{0}^{(i)})-\phi(s_{0}^{(i)})\right)^{2},$ (14)
where $N$ is the number of trajectories in the batch.
In addition to the MSE loss, we incorporate auxiliary loss terms to enforce the Lyapunov properties:
  • Positive Definiteness: We enforce $V(s)>0$ for all $s\neq 0$ and $V(0)=0$.
  • Decrease Condition: To promote stability, we penalize cases where the Lyapunov function does not decrease along the trajectory. For instance, using a hinge loss:
    $\mathcal{L}_{\text{dec}}=\frac{1}{N}\sum_{i=1}^{N}\max\{0,V_{\theta}(s_{t+1}^{(i)})-V_{\theta}(s_{t}^{(i)})+\epsilon\},$ (15)
    where $\epsilon>0$ is a margin.
The overall Lyapunov loss is then defined as a weighted combination of these terms:
$\mathcal{L}_{\text{Lyapunov}}=\mathcal{L}_{\text{MSE}}+\alpha\,\mathcal{L}_{\text{dec}},$ (16)
with $\alpha$ being a weighting factor chosen empirically. The learned Lyapunov function can be used to study the stability guarantees of a trained controller.
Finally, to encourage stability in the decision-making process, the learned neural Lyapunov function is incorporated into the loss of the Decision Transformer. The final loss for the Decision Transformer becomes:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DT}}+\gamma\,\mathcal{L}_{\text{Lyapunov}},$ (17)
where $\mathcal{L}_{\text{DT}}$ denotes the original Decision Transformer loss and $\gamma$ is a hyperparameter that balances the contribution of the Lyapunov loss. By minimizing $\mathcal{L}_{\text{total}}$, the Decision Transformer is guided to select actions that not only maximize reward but also lead the system towards more stable states as indicated by a decreasing $V_{\theta}(s)$.
@@ V-A Testing Cases and Implementation Details
We evaluate our method across several failure and damage scenarios for the UAV. Unlike the random and independent parameter modifications used during training, testing cases involve modifications to aerodynamic parameters and control surfaces based on the changes that a specific failure produces in the UAV’s dynamics, creating unseen situations for the agent. The defined scenarios, inspired by Refs. [10, 11], are:
@@
Experiments and training are conducted using JSBSim, a high-fidelity 6-DoF flight dynamics model. The UAV physics model used operates within flight conditions of $h\in[4000,24000]$ ft and $V_{T}\in[260,360]$ ft/s. Episodes have a maximum length of 1000 steps. To account for real-world uncertainties, we incorporate external perturbations and sensor noise during both training and testing. Wind gusts are applied following a Gaussian distribution N(0, 2 ft/s), introducing stochastic disturbances. Additionally, observation noise –modeled by a Gaussian distribution N(0, 0.02x)– is injected into sensor measurements to simulate real-world inaccuracies, ensuring the robustness of both the teacher and student models. These measures help bridge the simulation to real (sim-to-real) gap, improving generalization to real UAV conditions and paving the way for future real-world validation.
@@ V-B Comparison Study
Damaged Wing $\mu\pm\sigma$318.68 ± 20.46 -379.25 ± 91.68 -1394.93 ± 155.40 -1167.78 ± 338.98 -762.48 ± 444.57
Crash % 4.00% 9.00% 100.00% 92.00% 60.00%
@@
To assess the effectiveness of the DT under POMDP conditions, we compare the results of the DT student agent and the privileged teacher agent. Both policies are tested across all scenarios in 100 episodes each, with flight conditions and random seeds standardized for fair comparison. Fig. 5 shows the results, using the mean episode return as the performance metric, where the theoretical maximum is zero. Results demonstrate similar performance between the agents, showing the DT’s capability to learn behaviors from the expert dataset under the POMDP setting and to adapt to dynamic variations in context. The student model’s ability to match or outperform the teacher despite limited observations arises from the nature of offline reinforcement learning (RL). Rather than directly imitating the teacher, the student extracts essential decision patterns from expert trajectories, prioritizing key contextual cues for robust adaptation. This aligns with the findings in Ref. [31], where DTs demonstrate the ability to outperform their training data by using return-conditioning to generalize more effectively under uncertainty. Despite privileged information, the teacher may suffer from unstable training, whereas the student, trained on diverse replay data, learns a more adaptive strategy for handling unforeseen failures. Moreover, incorporating the Lyapunov-based loss term into the offline RL training further encourages the DT student to select actions that lead to more stable states, thereby improving overall return.
TABLE III: Student Model Architectures
Scenarios DT (Ours) LSTM TCN
Nominal $e_{h}$ 63.3 ± 55.4 102.3 ± 34.2 124.1 ± 41.7
$e_{\Psi}$ 2.4 ± 2.3 3.5 ± 2.2 3.9 ± 2.9
$e_{V_{T}}$ 4.6 ± 3.4 6.5 ± 3.7 9.1 ± 4.9
D. Wing $e_{h}$ 176.9 ± 90.6 1265.3 ± 1104.1 1823.6 ± 1331.9
$e_{\Psi}$ 12.8 ± 6.6 14.7 ± 5.3 16.4 ± 5.8
$e_{V_{T}}$ 9.4 ± 5.9 54.9 ± 40.5 69.3 ± 43.1
  • Tracking error performance comparison (in mean and std) between DT (Ours) and LSTM and TCN architectures. Errors are measured as L2 distance. $e_{h}$ in ft, $e_{\Psi}$ in degrees and $e_{V_{T}}$ in ft/s.
While transformers require higher computational resources than traditional sequence models, their parallelization may enable efficient inference, making them feasible for real-time UAV control in adequate hardware. Our model remains lightweight, ensuring real-time feasibility, as has been demonstrated in similar-size models in Ref. [41, 42, 43, 44]. Moreover, a trade-off between computational overhead and performance should be made. As shown in Table III, our DT achieves significantly lower tracking errors than LSTM and Temporal Convolutional Network (TCN), especially in failure scenarios. These results validate the superiority of transformer-based control in both accuracy and robustness.
Figure 6: Attention maps of the Decision Transformer (DT) before and after a failure occurs. Attention values are averaged across heads, layers, and $(o_{t},a_{t},r_{t})$ tuples over the 60-step context window. (a) Nominal case before failure, showing evenly distributed attention across past time steps. (b) Post-failure case (failure at timestep 100, current time at 140)
One key concern with transformer-based control systems in safety-critical applications is their black-box nature. To improve interpretability, we analyze the attention mechanism of the Decision Transformer (DT) to understand how it identifies failures and adapts its control policy. Figure 6 shows the attention patterns before and after a failure occurs, over the 60-step context window. In the nominal case (Figure 6(a)), attention is distributed evenly across past states. However, after a failure at timestep 100 (where the current timestep is 140), the DT adapts by focusing more on the time steps following the failure (Figure 6(b)), indicating a shift in control strategy based on recent dynamics. This adaptation mechanism enhances robustness by dynamically weighting relevant information in response to sudden changes in UAV behavior.
@@
Our method takes advantage of the in-context learning capabilities of transformers to adapt policy behavior during flight without changing weights [30], rather than modifying controller gains. Fig. 4 presents sample trajectories of our method versus the other algorithms in the comparison study for tracking $h$, $\Psi$, and $V_{T}$. In Fig. 4(a), the nominal UAV dynamics are shown, where all designed controllers track reference values with similarly high performance, with some noise in the measured airspeed due to light wind gusts.
@@
V-C Stability Analysis
Under partial observability and lacking an explicit model-based design, conventional methods for guaranteeing stability may be theoretically infeasible. However, following the findings in Refs. [20, 21, 22, 18, 24], we can use the learned Neural Lyapunov function to derive partial guarantees of stability and safety. Because the learned function satisfies the properties of a Lyapunov function, a controller that drives the system towards stability must exhibit a decreasing Lyapunov function, $V_{\theta}(s)$, throughout the episode, eventually reaching zero if the system is fully stabilized at
$s\in[\epsilon_{h},\epsilon_{\Psi},\epsilon_{V_{T}},\phi,p,q,r]=\hat{0}.$
To further examine the impact of including the Lyapunov loss term $\mathcal{L}_{\text{Lyapunov}}$ in Decision Transformer (DT) training, we evaluate the Lyapunov-based stability of both a baseline DT and the Lyapunov-enhanced DT. We conduct 100 experiments for each model under both nominal and failure scenarios (combining different types of failures).
Fig. 8 presents the mean and standard deviation of the Neural Lyapunov function, $V_{\theta}(s)$, over the timesteps in an episode—providing a measure of the “energy” in the system. If the Lyapunov energy tends to zero, it indicates that the system is converging to a stable equilibrium state. In Fig. 8(a), we observe that under nominal conditions, both DT models (trained with and without Lyapunov) display asymptotic stability, converging to minimal energy. The spike at timestep 200 arises from an intentional change in the reference values for altitude, heading, and airspeed ($h,\Psi,V_{T}$); after this disturbance, both models settle into a stable state.
Fig. 8(b) shows the failure scenario. A failure occurs at timestep 100, causing a surge in energy that both models stabilize quickly. However, once new reference values are introduced at timestep 200, the Lyapunov-enhanced DT converges to near-zero energy, while the baseline DT’s energy gradually increases until the episode ends. This behavior indicates that incorporating Lyapunov-based ideas into DT training can promote stability at inference.
Figure 8: Mean and standard deviation of the Neural Lyapunov function, $V_{\theta}(s)$, over time for the baseline DT and the Lyapunov-enhanced DT. (a) Nominal conditions, showing both models converging to minimal energy; (b) Failure scenario, illustrating how the Lyapunov-enhanced DT maintains lower energy, indicating greater stability.
V-D Discussion
While several nonlinear fault-tolerant control methods have been proposed—such as those based on fuzzy logic [7] and sliding mode control [8]—their adoption in real-world applications remains limited. One of the main reasons is that these approaches often require complex, computationally intensive routines for continuously monitoring system dynamics and adjusting control laws online. This computational overhead is a critical drawback for embedded platforms, where processing power and memory are tightly constrained.
In contrast, traditional cascade PID controllers, which are employed in widely used autopilot systems such as PX4 and Ardupilot, offer a much simpler and computationally efficient solution for low-level attitude control. Their hierarchical structure minimizes the need for online computations and parameter re-tuning, making them particularly well suited for microcontroller-based implementations. However, the simplicity of cascade PIDs also limits their performance when the UAV dynamics deviate significantly from the nominal model.
Our approach addresses these computational constraints by shifting the heavy computational burden to the offline training phase. The resulting Decision Transformer (DT) model, with only around 0.8M parameters, is designed for efficient real-time inference. Notably, recent studies have demonstrated that transformer models of similar size can be successfully deployed on microcontrollers—such as ARM Cortex-M devices—through aggressive quantization techniques that reduce both memory footprint and computational load [41, 42, 43, 44]. With quantization (e.g., converting weights and activations to 8-bit representations), the DT model’s inference can be optimized to meet the real-time constraints imposed by onboard flight controllers.
By eliminating the need for online re-parameterization, weight adaptation, or explicit fault identification during operation, our DT model circumvents the latency and computational complexity inherent in many nonlinear fault-tolerant control methods. This makes it a promising alternative for robust UAV control, effectively bridging the gap between advanced learning-based strategies and the stringent hardware limitations of embedded systems.